# WIMP/FIMP dark matter and primordial black holes with memory burden effect

### v3

@@ Abstract
The lifetime of primordial black holes (PBHs), which formed in the early universe, can be extended by the memory burden effect. Light PBHs may exist today and be candidates for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via the Hawking radiation of PBHs, and PBHs that survived Hawking evaporation via the memory burden effect. Feebly interacting massive particles (FIMPs) are alternatives to WIMPs. This paper shows that a small memory burden is preferable if the thermal production of WIMPs or FIMPs is much larger than the effect of PBH particle production via Hawking radiation. In addition, we show that the lower limit of the DM mass, called the warm DM (WDM) constraint, decreases with the memory burden effect of PBHs. Results suggest that the WDM constraint is more effective for FIMPs than for WIMPs.
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Section 3, we discuss the constraints on PBHs from BBN, cosmic microwave background radiation (CMB), GWs, and warm DM (WDM).We show that the lower limit of the DM mass (the WDM constraint, which is related to structure formation in the universe) decreases with the memory burden effect of PBHs. In Section 4, we assess the combined effects of PBHs and thermally produced WIMPs/FIMPs on the relic abundance of DM. Finally, Section 5 provides a summary of this study and its findings.
@@ 2.1 Time evolution
by solving $M_{\rm BH}(t_{\rm ev}^{0})=0$, where $t_{\rm ev}^{0}$ is the evaporation time.
@@
Approximation ($t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$) is adequate in some cases.
@@ 2.2 Initial density
The initial density of a PBH is
@@
As addressed in Ref. [25], $\beta<\beta_{\rm c}$ is a case of interest because we assume that PBHs will evaporate completely after the present time due to the memory burden effect and PBHs that have survived until the present day are DM candidates (no early-matter-dominated era is considered). Hereinafter, $\beta<\beta_{\rm c}$ should hold.
@@ 2.3 Survived PBH density
We assume that PBHs will evaporate completely after the present time due to the memory burden effect. Thus, the density of survived PBHs is [58]
@@
Using the relation
$\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}=\frac{q\rho_{\rm BH}^{\rm in}}{T_{\rm in}^{3}}=\frac{q\beta\rho_{r}^{\rm in}}{T_{\rm in}^{3}}$ (21)
and Eq. (19), we obtain (Eq. (3.37) in [25])
@@
2.4 Particle production
The number of particle species, $i$, from PBH evaporation is controlled by the relation between the particle mass, $m_{i}$, and the PBH formation temperature, which is
$T_{\rm in}=T_{\rm BH}(t_{\rm in})=\frac{M_{\rm P}^{2}}{M_{\rm in}}.$ (23)
If $m_{i}<T_{\rm in}$, then particle emission from a PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. If $m_{i}>T_{\rm in}$, then particle emission starts from time $t_{i}$, where $m_{i}=T_{\rm BH}(t_{i})$. The relic abundance of emitted particle DM is given by Eq. (3.33) in [25].
$\Omega_{\rm ev}h^{2}=1.8\times 10^{8}N_{\chi}\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\frac{m_{\chi}}{\rm GeV},$ (24)
where
$N_{\chi}=\frac{15\xi g_{\chi}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\times\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases},$ (25)
where $\xi=1$ for bosons and $\xi=3/4$ for fermions.
@@ 3.3 GWs
Figure 1: WDM constraint with memory burden effect of PBHs. Left panel: lower limit of WDM mass vs. initial mass of PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. Right panel: $q$ dependence on lower limit of $m_{\chi}$.
@@ 3.4 WDM
In cases with negligible DM interaction with any particles in the thermal bath, DM particles can only lose energy (or momentum) by redshifting. The present redshifted momentum of DM, $p(t_{0})$, is [48]
$p(t_{0})$ $=$ $\frac{a(t_{\rm ev})}{a(t_{0})}p(t_{\rm ev})$ (28)
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{a(t_{\rm eq})}{a(t_{0})}p(t_{\rm ev})$
$=$ $\frac{a(t_{\rm ev})}{a(t_{\rm eq})}\frac{\Omega_{\rm r}}{\Omega_{\rm m}}p(t_{\rm ev}),$
where $\Omega_{\rm r}$ and $\Omega_{\rm m}$ are the density parameters of radiation and matter, respectively. Using $\rho_{\rm r}\propto a^{-4}$, we obtain
$\frac{a(t_{\rm ev})}{a(t_{\rm eq})}=\left(\frac{\rho_{\rm r}(T_{\rm eq})}{\rho_{\rm r}(T_{\rm ev})}\right)^{1/4}.$ (29)
We can use the relations
$\rho_{\rm r}(T_{\rm eq})$ $=$ $\rho_{\rm c}\left(\frac{a(T_{0})}{a(T_{\rm eq})}\right)^{3}=\rho_{\rm c}a(T_{\rm eq})^{-3},$
$\rho_{\rm r}(T_{\rm ev})$ $=$ $3M_{\rm P}^{2}H^{2}=\frac{3M_{\rm P}^{2}}{(2t_{\rm ev})^{2}}\simeq\frac{3M_{\rm P}^{2}}{4\tau^{2}},$
$a(T_{\rm eq})$ $=$ $a(t_{\rm eq})=\frac{\Omega_{\rm r}}{\Omega_{\rm m}},$ (30)
and $p(t_{\rm ev})\simeq\langle E\rangle\simeq 6T_{\rm in}$ for light DM ($m_{\rm DM}\ll T_{\rm in}$), where $\rho_{\rm c}=3H_{0}^{2}/(8\pi G)=1.05h^{2}\times 10^{-5}\ {\rm GeV\ cm^{-3}}$ is the critical density of the universe and $\tau$ is the PBH lifetime.
Without the memory burden effect, we set $\tau=\tau^{0}$ under the assumption of $t_{\rm in}\ll t^{0}_{\rm ev}$. In this case, the momentum of a DM particle today is
$p(t_{0})=1.53\times 10^{-12}\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{1/2}~{\rm GeV}$ (31)
at $g_{*}(T_{\rm BH})=106.75$, $\Omega_{\rm r}=5.38\times 10^{-5}$, and $\Omega_{\rm m}=0.315$ [107]. Assuming the WDM is heavier than 3.5 keV [108], the upper limit of the DM velocity is approximately $v_{\rm DM}\lesssim 1.8\times 10^{-8}$ [109]. From $p(t_{0})=m_{\rm DM}v_{\rm DM}$, we obtain the following WDM constraint on the DM mass and the initial mass of the PBH:
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (32)
Considering the memory burden effect, we can utilize the final time for the semiclassical era, $t_{\rm q}$, as an effective lifetime of PBHs for particle production. In this case, we set $\tau=\tau^{0}(1-q^{3})$ instead of $\tau=\tau^{0}$ from Eq. (10). The WDM constraint with the memory burden effect is
$m_{\rm DM}\gtrsim 4.07\times 10^{-2}(1-q^{3})^{1/2}\left(\frac{M_{\rm in}}{1{\rm g}}\right)^{1/2}~{\rm GeV}.$ (33)
Thus, the allowed minimum DM mass is reduced by the memory burden effect.
Figure 1 shows the nature of Eq.(33). The left panel shows the lower limit of the WDM mass against the initial mass of the PBH. The left (right) gray region is excluded from the CMB (BBN) constraint. The blue solid line indicates the prediction without the memory burden effect, whereas the dotted lines show predictions with the memory burden effect at q=0.0050.9q=0.005-0.9. The right panel shows $q$ dependence on the lower limit of $m_{\chi}$. As expected, a large memory burden effect of the PBH reduces the WDM constraint.

### v4

@@ Abstract
The lifetime of primordial black holes (PBHs), which formed in the early universe, can be extended by the memory burden effect. Light PBHs may exist today and be candidates for dark matter (DM). We assume that DM is made of thermally produced weakly interacting massive particles (WIMPs), WIMPs produced via the Hawking radiation of PBHs, and PBHs that survived Hawking evaporation via the memory burden effect. Feebly interacting massive particles (FIMPs) are alternatives to WIMPs. Focusing on parameter regions where thermal production dominates and PBHs never dominate the energy density of the Universe, we identify a sufficient condition under which DM particles emitted from PBHs do not thermalize with the thermal bath. In this regime, the total DM relic abundance can be consistently obtained as the sum of the three components. In addition, we show that the contribution from gravitational freeze-in via graviton exchange remains subdominant within the parameter space considered.
@@ 1 Introduction
The remainder of this paper is organized as follows. Section 2 provides a brief review of PBHs with the memory burden effect. In Section 3, we discuss the constraints on PBHs from BBN, cosmic microwave background radiation (CMB), GWs, and warm DM (WDM). In Section 4, we assess the combined effects of PBHs and thermally produced WIMPs/FIMPs on the relic abundance of DM. Finally, Section 5 provides a summary of this study and its findings.
@@ 2.1 Time evolution
by solving $M_{\rm BH}(t_{\rm ev}^{0})=0$, where $t_{\rm ev}^{0}$ is the evaporation time. The PBHs with $M_{\rm in}\lesssim 10^{9}$ g have evaporated before BBN, we use $t_{\rm BBN}=1$ s, whereas PBHs with $M_{\rm in}\gtrsim 10^{15}$ g can survive until today without memory burden effect.
@@
Approximation ($t_{\rm ev}^{k}\simeq 1/\Gamma_{\rm BH}^{k}$) is adequate in some cases. For example, for $k=1$ and $q=1/2$, the PBH with $M_{\rm in}\gtrsim 2\times 10^{7}$ g can survive until today due to the memory burden effect.
@@ 2.2 Initial density
The initial density of a PBH is usually characterized by the dimensionless parameter
@@
2.3 Particle production
An origin of the DM is an evaporating PBHs. According to Ref. [25], we divide the era of Hawking radiation into two phases as follows:
Phase-I:
Semiclassical evaporation phase.
Phase-II:
The second phase where quantum correction is effective.
Let $N_{1i}$ and $N_{2i}$ be the number of particles which are emitted from PBH in phase-I and phase-II, respectively. The emission rate of particle species $i$ with mass $m_{i}$ and internal degrees of freedom $g_{i}$ per unit time per unit energy due to Hawking radiation is given by [25]
$\frac{d^{2}N_{1i}}{dEdt}$ $=\frac{27}{4}\frac{g_{i}}{32\pi^{3}}\frac{(E/T_{\rm BH})^{2}}{\exp(E/T_{\rm BH})\pm 1},$ (20)
$\frac{d^{2}N_{2i}}{dEdt}$ $=\frac{1}{\left[S(M_{\rm BH})\right]^{k}}\frac{d^{2}N_{1i}}{dEdt},$ (21)
where the sign $\pm$ is used for fermions and bosons, respectively. After integrating for all energy mode, we obtain
$\frac{dN_{1i}}{dt}$ $=\frac{27}{4}\frac{\xi g_{i}\zeta(3)}{16\pi^{3}}\frac{M_{\rm P}^{2}}{M_{\rm BH}},$ (22)
$\frac{dN_{2i}}{dt}$ $=\frac{27}{4}\frac{\xi g_{i}\zeta(3)2^{k}}{16\pi^{3}}\frac{M_{\rm P}^{2+2k}}{M_{\rm BH}^{1+2k}},$ (23)
where $\xi=1$ for bosons and $\xi=3/4$ for fermions.
The total number of particle species $i$ from evaporation of BH is controlled by the relation between the particle mass $m_{i}$ and the PBH formation temperature
$T_{\rm in}=T_{\rm BH}(t_{\rm in})=\frac{M_{\rm P}^{2}}{M_{\rm in}}.$ (24)
If $m_{i}<T_{\rm in}$, the particle emission from PBH occurs from $t_{\rm in}$ to $t_{\rm ev}$. On the other hand, if $m_{i}>T_{\rm in}$, the particle emission will start from a time where $m_{i}=T_{\rm BH}(t_{i})$. From Eqs. (4), (11), and (12), we obtain
[…]
with assumption of $t_{i}\gg t_{\rm in}$.
For $m_{i}<T_{\rm in}$, integrating Eqs. (22) within $t_{\rm in},t_{\rm q}]$ and (23) within $t_{\rm q},t]$, the number of particle species $i$ is given by
$N_{1i}$ $=\frac{15\xi g_{i}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}(1-q^{2})\frac{M_{\rm in}^{2}}{M_{\rm P}^{2}},$ (26)
$N_{2i}$ $=\frac{15\xi g_{i}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\frac{q^{2}M_{\rm in}^{2}}{M_{\rm P}^{2}}\left[1-\left(1-\frac{t-t_{\rm q}}{t_{\rm ev}}\right)^{2/(3+2k)}\right].$ (27)
On the other hand, there are two cases for $m_{i}>T_{\rm in}$: $t_{i}<t_{\rm q}$ and $t_{i}>t_{\rm q}$. For $t_{i}<t_{\rm q}$, integrating Eqs. (22) within $t_{i},t_{\rm q}]$ and (23) within $t_{\rm q},t]$, the number of particle species $i$ is given by
$N_{1i}$ $=\frac{15\xi g_{i}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\left(\frac{M_{\rm P}^{2}}{m_{i}^{2}}-\frac{q^{2}M_{\rm in}^{2}}{M_{\rm P}^{2}}\right),$ (28)
$N_{2i}$ $=\frac{15\xi g_{i}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\frac{q^{2}M_{\rm in}^{2}}{M_{\rm P}^{2}}\left[1-\left(1-\frac{t-t_{\rm q}}{t_{\rm ev}}\right)^{2/(3+2k)}\right].$ (29)
For $t_{i}>t_{\rm q}$, corresponding to $m_{i}>M_{\rm P}^{2}/(qM_{\rm in})=T_{\rm in}/q$, we obtain
$N_{1i}$ $=0,$ (30)
$N_{2i}$ $=\frac{15\xi g_{i}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\frac{q^{2}M_{\rm in}^{2}}{M_{\rm P}^{2}}\left[\left(1-\frac{t_{i}-t_{\rm q}}{t_{\rm ev}}\right)^{2/(3+2k)}-\left(1-\frac{t-t_{\rm q}}{t_{\rm ev}}\right)^{2/(3+2k)}\right].$ (31)
2.3.1 $\beta>\beta_{\rm c}$
For $\beta>\beta_{\rm c}$, PBHs dominate the Universe before PBH decays. Since a PBH behaves as matter ($H=2/(3t)$) and according to the Friedmann equation, we obtain
$H^{2}(a_{\rm ev})=\frac{\rho_{r}(a_{\rm ev})}{3M_{\rm P}^{2}}=\frac{1}{3M_{\rm P}^{2}}\frac{\pi^{2}g_{\ast}T^{4}(a_{\rm ev})}{30}=\frac{4(\Gamma_{\rm BH}^{k})^{2}}{9},$ (32)
and
$T(a_{\rm ev})=M_{\rm P}\left(\frac{4}{3}\frac{30}{\pi^{2}g_{\ast}}\right)^{1/4}\left[(3+2k)2^{k}\epsilon\left(\frac{M_{\rm P}}{qM_{\rm in}}\right)^{3+2k}\right]^{1/2}.$ (33)
In the PBH dominated era, PBH energy density gets converted into radiation density entirely through the evaporation. The reheating temperature would be the temperature of the decay products [25]. If we suppose an instant thermalization, we can use the relation of $\rho_{\rm BH}(a_{\rm ev})=\rho_{r}(a_{\rm ev})\simeq n_{\rm BH}(a_{\rm ev})qM_{\rm in}$. We have
$n_{\rm BH}(a_{\rm ev})=\frac{4}{3}M_{\rm P}^{3}(3+2k)^{2}2^{2k}\epsilon^{2}\left(\frac{M_{\rm P}}{qM_{\rm in}}\right)^{7+4k}.$ (34)
The relic abundance of particle DM $\chi$ from PBH evaporation is obtained as [58]
$\Omega_{\chi}^{\rm ev}h^{2}=1.6\times 10^{8}\frac{g_{0}}{g_{\rm ev}}\frac{n_{\chi}(a_{\rm ev})}{T^{3}(a_{\rm ev})}\frac{m_{\chi}}{\rm GeV},$ (35)
where $g_{0}=3.91$ and $g_{\rm ev}\simeq 106.75$ are the effective number of degrees for entropy density at today and at the end of evaporation, respectively. $n_{\chi}(a_{\rm ev})=n_{\rm BH}(a_{\rm ev})N_{\chi}$ is the number density of particle DM at the end of evaporation. The relic abundance of particle DM emitted by Hawking radiation is given by (Eqs.(3.22) and (3.23) in Ref.[25])
$\Omega_{\chi}^{\rm ev}h^{2}=1.9\times 10^{7}N_{\chi}g^{1/2}_{\ast}(T_{\rm BH})\left(2^{k}(3+2k)\right)^{1/2}\left(\frac{M_{\rm P}}{qM_{\rm in}}\right)^{5/2+k}\frac{m_{\chi}}{\rm GeV},$ (36)
where
$N_{\chi}=\frac{15\xi g_{\chi}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\times\begin{cases}\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2},&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases}.$ (37)
We assume $g_{\ast}(T_{\rm BH})\simeq 106.75$. For $k\rightarrow 0$ and $q\rightarrow 1$, the expression of $\Omega_{\chi}^{\rm ev}h^{2}$ reduces to the standard semi-classical expression [48, 49].
2.3.2 $\beta<\beta_{\rm c}$
For $\beta<\beta_{\rm c}$, we suppose that the standard Hawking evaporation (phase-I) occurs before BBN and the PBHs evaporate completely after the present time due to the memory burden effect (phase-II) citeHaque2024JCAP.
The Hubble parameter at the time of evaporation is given by
$H^{2}(a_{\rm ev})=\frac{\rho_{r}(a_{\rm ev})}{3M_{\rm P}^{2}}=\frac{1}{3M_{\rm P}^{2}}\frac{\pi^{2}g_{\ast}T^{4}(a_{\rm ev})}{30}=\frac{(\Gamma_{\rm BH}^{k})^{2}}{4}.$ (38)
From this relation, we obtain the evaporation temperature as
$T(a_{\rm ev})=M_{\rm P}\left(\frac{3}{4}\frac{30}{\pi^{2}g_{\ast}}\right)^{1/4}\left[(3+2k)2^{k}\epsilon\left(\frac{M_{\rm P}}{qM_{\rm in}}\right)^{3+2k}\right]^{1/2}.$ (39)
The relic abundance of particle DM $\chi$ from PBH evaporation is obtained as [58]
$\Omega_{\chi}^{\rm ev}h^{2}=1.6\times 10^{8}\frac{g_{0}}{g_{\rm q}}\frac{n_{\chi}(a_{\rm q})}{T^{3}(a_{\rm q})}\frac{m_{\chi}}{\rm GeV},$ (40)
where $g_{\rm q}\simeq 106.75$ is the effective number of degrees for entropy density at the end of phase-I. $n_{\chi}(a_{\rm q})$ is the number density of particle DM at the end of phase-I. With the relation of
$\frac{n_{\chi}(a_{\rm q})}{T^{3}(a_{\rm q})}=\frac{N_{1\chi}n_{\rm BH}(a_{\rm in})}{T_{\rm in}^{3}}=\frac{N_{1\chi}\rho_{\rm BH}^{\rm in}}{T_{\rm in}^{3}M_{\rm in}}=\frac{N_{1\chi}\beta\rho_{r}^{\rm in}}{T_{\rm in}^{3}M_{\rm in}},$ (41)
and Eq.(19), the relic abundance of particle DM emitted in the semiclassical phase (phase-I) is given by (Eq.(3.33) in Ref.[25])
$\Omega_{\chi}^{\rm ev}h^{2}=1.8\times 10^{8}N_{\chi}\beta\left(\frac{M_{\rm P}}{M_{\rm in}}\right)^{3/2}\frac{m_{\chi}}{\rm GeV},$ (42)
where
$N_{\chi}=\frac{15\xi g_{\chi}\zeta(3)}{g_{\ast}(T_{\rm BH})\pi^{4}}\times\begin{cases}(1-q^{2})\left(\frac{M_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}<T_{\rm in}\\ \left(\frac{M_{\rm P}}{m_{\chi}}\right)^{2}-\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{2},&{\rm for}\ m_{\chi}>T_{\rm in}\\ \end{cases}.$ (43)
For $q\rightarrow 0$, the expression of $N_{\chi}$ reduces to the standard semi-classical expression for $m_{\chi}<T_{\rm in}$ and $m_{\chi}>T_{\rm in}$ [48, 49].
@@ 2.4 Survived PBH density
According to Ref.[25], the $\beta<\beta_{c}$ case is interesting because we assume that PBHs evaporate completely after the present time due to memory burden effect and the survived PBHs until today are also a candidate of DM (there is no “early” matter dominated era). Hereafter, we use the relations for $\beta<\beta_{c}$.
If PBHs evaporate completely after the present time due to the memory burden effect (phase-II). The survived PBHs until today are also a candidate of DM. The density of survived PBHs is obtained as [58]
@@
With the relation of
$\frac{\rho_{\rm BH}(a_{\rm q})}{T^{3}(a_{\rm q})}=\frac{q\rho_{\rm BH}^{\rm in}}{T_{\rm in}^{3}}=\frac{q\beta\rho_{r}^{\rm in}}{T_{\rm in}^{3}},$ (45)
and Eq.(19), we obtain (Eq. (3.37) in Ref.[25])
@@ 3.4 WDM
According to Refs. [109, 112, 30], DM particles produced via PBH evaporation must satisfy a lower bound on their mass,
$m_{\rm DM}\gtrsim 10^{4}\langle E_{\rm DM}(t_{\rm eq})\rangle,$ (49)
where $\langle E_{\rm DM}(t_{\rm eq})\rangle$ denotes the average kinetic energy of DM particles at the epoch of radiation-matter equality. The average kinetic energy at this epoch can be estimated by redshifting the energy at the evaporation time,
$\langle E_{\rm DM}(t_{\rm eq})\rangle\simeq\langle E_{\rm DM}(t_{\rm ev})\rangle\frac{a_{\rm ev}}{a_{\rm eq}}\simeq\delta T_{\rm BH}\frac{T(a_{\rm eq})}{T(a_{\rm ev})}\left[\frac{g_{\ast s}(T(a_{\rm eq}))}{g_{\ast s}(T(a_{\rm ev}))}\right]^{1/3}$ (50)
Here $T_{\rm BH}\simeq M^{2}_{\rm P}/(qM_{\rm in})$, and $T(a_{\rm eq})\simeq 0.75~{\rm eV}$ is the radiation temperature at radiation-matter equality. The function $g_{\ast s}(T)$ represents the effective number of relativistic degrees of freedom contributing to the entropy density.
Following Refs.[113, 109], we introduce a numerical factor $\delta\simeq 1.3$ to parametrize the average DM energy at evaporation as $\langle E_{\rm DM}(t_{\rm ev})\rangle\simeq\delta T_{\rm BH}$. We also assume entropy conservation from the PBH evaporation epoch until radiation-matter equality.
According Eqs. (33) and (39), the lower bound on the DM mass is given by
$m_{\rm DM}$ $\gtrsim$ $g_{\ast}^{1/4}\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{\frac{1+2k}{2}}\frac{1}{\sqrt{(3+2k)2^{k}\epsilon}}\left[\frac{g_{\ast s}(T(a_{\rm eq}))}{g_{\ast s}(T(a_{\rm ev}))}\right]^{1/3}$ (51)
[…]
$\gtrsim$ $\left(\frac{qM_{\rm in}}{M_{\rm P}}\right)^{\frac{1+2k}{2}}\frac{1}{\sqrt{(3+2k)2^{k}}}\times\begin{cases}3.39\times 10^{-6}~{\rm GeV}&{\rm for}\ \beta>\beta_{\rm c}\\ 3.88\times 10^{-6}~{\rm GeV}&{\rm for}\ \beta<\beta_{\rm c}\\ \end{cases},$
where in the last line, we set $g_{\ast s}(T(a_{\rm eq}))=3.91$ and $g_{\ast}=g_{\ast s}(T(a_{\rm ev}))\simeq 106.75$. For example, we obtain
$m_{\rm DM}\gtrsim\left(\frac{q}{0.5}\right)\left(\frac{M_{\rm in}}{10~{\rm g}}\right)\times\begin{cases}1.6~{\rm GeV}&{\rm for}\ \beta>\beta_{\rm c}\\ 1.9~{\rm GeV}&{\rm for}\ \beta<\beta_{\rm c}\\ \end{cases},$ (52)
for $k=1/2$. For a given PBH mass, this bound becomes stronger for larger values of $k$ and $q$ , since in that case PBH evaporation is delayed; as a result, the average kinetic energy of the radiated particles at radiation-matter equality increases [30].